I was wondering about doing Slope Intercept form and Standard form. When you are converting from Slope Intercept form to Standard form, and there are two fractions, what do you do?
For example: Slope Intercept form--> y= 2/3x + 5/9
I know you multiply everything by the denominator or in this case the least common denominator. However, I've checked the answer and it said it was 9y = 6x + 5. Wouldn't the least common denominator be 3?

Answer by richwmiller(17219)

y= 2/3x + 5/9
multiply by 9
9y=9*2/3 x+9*5/9
9y=18/3 x+5
9y=6x+5